RD 1801
Hobart 26m
Disk VSN: HOB+0131 | Data volume at beginning: 9101.6 GB
- 1800UT Started OK, 40k at 400K :( pressure is dropping and temperatures rising keeping an eye out (JS)
- 1930UT Helium pressure is above 5 umHg, can't do anything until Warren gets to MtP (JS)
- 23:00 stopped for an hour to change cryo hose. Still a leak (Warren)
- 00:00 receiver cool but helium requires constant topping up Warren)
- 00:00 data down about 170 GB due to service hour (warren)
- 10.42 UT, Cryogenics He supply pressure is warming up now (250 psi). 70 K stage temperature around 125.9 K and pressure 7.1 micorns Hg. Have informed Warren about it, he said just need to wait till experiment finish (mas)